OpenStudy (anonymous):

what is the derivative of ln(1+1/x)? they say the answer is xsquared-1/ xcubed+x now can someone explain how you find the answer

OpenStudy (anonymous):

the answer given is correct but not simplified.... use the formula \(\large [lny]'=\frac{y'}{y} \) to find the derivative...

OpenStudy (anonymous):

then how do i use this formula for the problem

OpenStudy (anonymous):

let \(\large y=1+\frac{1}{x} \) so, \(\large [ln(1+\frac{1}{x})]'=\frac{[1+\frac{1}{x}]'}{1+\frac{1}{x}} \) can you find \(\large [1+\frac{1}{x}]'= \) ????

OpenStudy (anonymous):

no how do u find 1+1/x? is the answer -1x-1/2

OpenStudy (anonymous):

use the sum rule for derivatives... \(\large [1+\frac{1}{x}]'=[1]'+[\frac{1}{x}]' \) continue....

OpenStudy (anonymous):

 ??

OpenStudy (anonymous):

ok... good... so, \(\large [ln(1+\frac{1}{x})]'=\frac{[1+\frac{1}{x}]'}{1+\frac{1}{x}}=\frac{-x^{(\frac{-1}{2})}}{1+\frac{1}{x}} \) now just simplify...

OpenStudy (anonymous):

wait but how do i simplify to get ...